Working alone, Carlos can oil the lanes in a bowling alley in five hours. Jenny can oil the same lanes in nine hours. If they work together how long would it take them?
Answer by macston(5194)

Carlos oils all lanes/5 hrs= 1/5 of the lanes/hr.
Jenny oils all lanes/9 hrs=1/9 lanes per hour.
Together they oil 1/5 per hr+1/9 per hr=9/45 per hr+ 5/45 per hr=14/45 lanes per hr=0.311=31.1% of all lanes per hour
To do all lanes 1/0.311=3.21 hrs ANSWER: Together it takes Carlos and Jenny 3.21 hours to oil all the lanes.